# On the Use of Large Language Models in Mutation Testing

### v2

@@ 3.5.3. Settings for RQ4
    Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the examples.
@@
3.5.4. Settings for RQ5
Generating non-compilable mutations incurs additional meaningless economic and time costs, so in the third part of our study, we study the root causes behind LLM generating non-compilable mutations. Specifically, we sample 384 non-compilable mutations of each mutation generation approach and analyze their reason for being rejected by the Java compiler and the features of their surrounding code context.
4. Evaluation Results
We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
4.1. RQ1: Performance on Cost and Usability
4.1.1. Cost
For each technique, we record the generation duration, as shown in the Avg. Gen. Time row of Table 4. In addition, we report the mutation count and mutation score, although they do not fully reflect the overall quality of the mutation set.
From the table, we find that GPT-3.5, GPT-4o, GPT-4o-Mini, DeepSeek, StarChat, and CodeLlama generate 538338, 539715, 588803, 594711, 463248, and 457195 mutations, respectively222 Although the LLMs are instructed to generate the same number of mutations, they do not consistently follow the instructions, thereby yielding different numbers of mutations., with a mutation score ranging from 0.516 to 0.726. Overall, traditional approaches (e.g., LEAM and PIT) yield lower mutation scores among all the approaches.
In terms of time cost, the six LLMs cost 1.79s, 1.76s, 1.65s, 4.25s, 7.53s, and 9.06s to generate mutations, respectively. In comparison, LEAM, $\mu$Bert, PIT, and Major take 3.06s, 2.23s, 0.02s, and 0.08s, respectively. The rule-based approaches, PIT and Major, are the fastest, while LLMs show a competitive speed with small model-based approaches, LEAM and $\mu$Bert.
Finding 1 (RQ1): Rule-based approaches cost less time than LLMs in generating mutations. For example, GPT LLMs take fewer than 1.8s per mutation. In contrast, PIT and Major only require 0.02s and 0.08s, respectively, while small model-based approaches (i.e., LEAM and $\mu$Bert) take 2-3s.
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
Mut. Score 0.697 0.726 0.708 0.694 0.516 0.700 0.545 0.675 0.581 0.486
Mut. Count 538338 539715 588803 594711 463248 457195 890417 200896 2890433 346282
Avg. Gen. Time 1.79 1.76 1.65 4.25 7.53 9.06 3.06 2.33 0.02 0.08
Comp. Rate 60.2% 75.6% 73.6% 75.5% 11.1% 70.2% 35.0% 22.5% —— 98.6%
Useless Mut. Rate 10.9% 7.8% 6.7% 8.3% 8.9% 39.3% 1.0% 1.7% 0.0% 0.0%
Eq. Mut. Rate 2.2% 1.2% 1.7% 1.2% 1.8% 1.2% 1.3% 2.5% 0.0% 0.6%
Real Bug Detec. 91.7% 93.4% 93.4% 92.8% 47.1% 83.1% 71.7% 71.3% 51.3% 74.4%
Coupling Rate 41.4% 43.6% 44.1% 41.5% 29.4% 39.8% 28.7% 41.6% 14.0% 36.0%
Ochiai Coeff. 65.0% 68.5% 66.9% 61.8% 19.3% 40.8% 37.4% 31.6% 31.5% 44.4%
4.1.2. Mutation Usability
The usability of mutations (i.e., whether the mutations can be used to calculate mutation score) is affected by the capabilities of LLMs in correctly understanding the task and generating compilable, non-duplicate, and non-equivalent mutations. The results are shown as the Comp. Rate, Useless Mut. Rate, and Eq. Mut. Rate rows in Table 4. Since the mutation generation process of the PIT model is not visible and it only outputs mutations that pass the tests, we excluded it from the compilability rate calculation.
Compilability Rate: Four out of the six LLMs (i.e., GPT-4o, GPT-4o-Mini, DeepSeek, CodeLlama) achieve over 70%, with one exceeding 60% (i.e., GPT-3.5). In contrast, the small model-based approaches, LEAM and $\mu$Bert, have rates at 35.0% and 22.5%, respectively. Compared to small models, LLMs like GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 40% higher than that of LEAM and $\mu$Bert. The rule-based approach Major significantly outperforms, with nearly 99% of its mutations successfully compiling. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true).
Useless Mutation Rate: We also observe that LLMs generate a significant portion of duplicate and useless mutations. For useless mutations, the percentage generated by LLMs ranges from a minimum of 6.7% (GPT-4o-Mini) to a maximum of 39.3% (CodeLlama), and five out of the six LLMs have a useless mutation rate below 11%. In contrast, all the traditional approaches rarely generate duplicate mutations, which is likely due to them being governed by strict grammar rules that limit the production of redundant mutations. For example, PIT and Major iteratively apply different mutation operators producing no redundant mutants.
Table 5. Sample Results for Equivalent Mutations
Dataset Type GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
D4J #All 537109 538509 587583 593481 462042 455634 888452 199347 2888641 345292
#Sampled 384 384 384 384 384 384 384 384 384 384
#Eq. Mut 9 3 5 4 8 5 6 8 0 2
CD #All 1229 1206 1220 1230 1206 1561 1965 1549 1792 990
#Sampled 293 292 293 293 292 309 322 308 317 277
#Eq. Mut 6 5 7 4 4 3 3 9 0 2
Equivalent Mutation Rate: Equivalent mutations change the source code syntactically without affecting the program’s functionality, thereby affecting the accuracy of mutation score calculations. Due to the undecidable nature of identifying equivalent mutations (Budd and Angluin, 1982), it is not feasible to design an algorithm that can completely detect them. Therefore, we sample a subset of mutations for each technique and manually assess whether they are equivalent. Following the common practice of sampling (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015), to maintain a 95% confidence level and a 5% margin of error, we sampled a fixed number of mutations from each approach’s output, to calculate the Equivalent Mutation Rate. Following the protocol of existing studies (Islam et al., 2019; Guo et al., 2024), two authors, each with over five years of Java programming experience, first received training on the equivalent mutants, and then independently labeled the mutations. The final Cohen’s Kappa coefficient score (Viera et al., 2005) is more than 90%, indicating a high level of agreement. All labeling results are available in our package for reproducibility (our, 2024), and Table 5 presents the sampling and labeling results.
As shown in the Eq. Mut. Rate row of Table 4, we again observe that the rule-based approaches, PIT and Major, perform the best in generating non-equivalent mutants, with an equivalent mutation rate of less than 1.0%. All six LLMs exhibit an equivalent mutation rate ranging from 1.2% to 2.2%, similar to LEAM and $\mu$Bert. Overall, the equivalent mutation rates for all approaches studied are relatively low and can be ignored during mutation testing. We further analyzed the mutations generated by PIT and Major, and found that they apply specific mutation operators to a statement (e.g., for a < b, only generating !(a < b), true, and false), significantly reducing the equivalent mutation rate.
Finding 2 (RQ1): Rule-based approaches, PIT and Major, outperform others in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, GPT-4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%. In contrast, Major excels with a compilability rate of 98.3%, a useless mutation rate of 0%, and an equivalent mutation rate of 0.6%.
Figure 2. Number of newly introduced AST nodes of the studied approaches and the few-shot examples.
4.1.3. AST Node Dirversity
To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 shows the number of introduced types, which reveals that LLM-based mutation approaches introduce more types of new AST nodes than other approaches. The best of LLMs, GPT-4o-Mini and DeepSeek, exhibit the greatest diversity, introducing 57 new AST node types. Among the traditional approaches, LEAM, $\mu$Bert, and Major introduce 31, 30, and 2 new types of AST nodes, respectively. Additionally, all six LLMs significantly are higher than traditional approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
Table 6. Ratio of Deletion Mutations and the Top-3 Newly Introduced AST Node Types for Each Approach.
Top K GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major
Del. 5.7% 2.2% 0.5% 1.1% 0.2% 0.1% 48.4% 0.6% 20.2%
1st LT(36.1%) LT(43.5%) LT(41.7%) LT(40.1%) LT(20.1%) LT(42.5%) LT(21.4%) LT(25.4%) ST(58.6%)
2nd BO(17.4%) BO(16.5%) BO(22.0%) BO(18.7%) MR(13.7%) BO(23.8%) MR(14.7%) SE(21.7%) LT(41.4%)
3rd MI(10.9%) MI(7.9%) MI(7.1%) MI(10.5%) BO(11.8%) MI(8.0%) MI(12.0%) MR(21.9%) ——
LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 6, which highlights the top three most frequent AST node types. For the proportion of deletions, LEAM shows the highest deletion mutation ratio, with 48.4%, followed by Major at 20.2%. In contrast, all six LLMs generate deletion mutations at a rate below 6%, with three models producing fewer than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we found that deletions are one of the key factors contributing to the lower compilation success rate of traditional approaches.
Regarding AST node types, the most frequent new AST node type of all the learning-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the rule-based approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
Finding 3 (RQ1): LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches, while less inclined to generation deletion mutations.
4.2. RQ2: Behavior Similarity
The bottom three rows of Table 4 present a comparative evaluation of Behavior Metrics for mutation generation approaches.
4.2.1. Real Bug Detectability
Among the six LLMs, five models (excluding StarChat) outperform all the traditional approaches. Of the traditional approaches, the rule-based approach Major performs the best, detecting 74.4% of the bugs. GPT-4o and GPT-4o-Mini achieve the highest detection rates, identifying 93.4% of the 851 bugs, excelling Major by 19% on all bugs.
In particular, on the ConDefects dataset, which mitigates data leakage risks, the real bug detection rates for the various approaches are as follows: GPT-3.5 (78.9%), GPT-4o (84.6%), GPT-4o Mini (82.5%), DeepSeek (83.7%), StarChat (32.9%), CodeLlama (63.4%), LEAM (48.9%), $\mu$Bert (47.2%), PIT (49.2%), and Major (56.1%). The best of LLMs (i.e., GPT-4o) outperforms nearly 29% over the best of traditional approaches (i.e., Major), highlighting the significant advantage that LLMs hold over traditional techniques in detecting real bugs.
Finding 4 (RQ2): In Real Bug Detectability, five out of six LLMs significantly outperform traditional approaches. In particular, on ConDefects, GPT-4o outperforms nearly 29% over the best of the traditional approach, Major, highlighting the advantages of LLMs in detecting real bugs.
4.2.2. Coupling Rate
The Coupling Rate measures the degree of coupling between the generated mutations and their corresponding real bugs.
Among traditional approaches, $\mu$Bert is the best one with a rate of 41.6%. However, two LLMs outperform all traditional approaches in this metric, achieved by the GPT-4 LLMs. The best LLM, GPT-4o-Mini, exhibits a rate of 44.1%, exceeding $\mu$Bert by 2.5%.
Finding 5 (RQ2): GPT-4 achieves the highest Coupling Rate at 44.1%, outperforming all traditional approaches, including the best traditional approach, $\mu$Bert, by 2.5%.
4.2.3. Ochiai Coefficient
The Ochiai Coefficient measures the semantic similarity between mutations and real bugs (Ojdanic et al., 2023a).
Similar to real bug detectability, most LLMs significantly outperform all existing approaches. Among the LLMs, GPT-4o achieves the highest Ochiai coefficient at 68.5%, 24.1% higher than the best rule-based approach, Major (44.4%), and 31.1% higher than the best small model-based approach, LEAM (37.4%). Moreover, five out of the six LLMs (excluding StarChat), outperforms all traditional approaches.
Finding 6 (RQ2): In the Ochiai Coefficient, five LLMs surpass all traditional generation approaches. In particular, the best LLM, GPT-4o, exceeds the best of traditional approaches, Major, by 24.1%.
4.3. RQ3: Impacts of Different LLMs
As shown in Table 4, we compare the six LLMs we studied in mutation generation. The results reveal significant performance differences among the models.
For Cost, GPT-4o-Mini is the fastest, requiring only 39% of the time taken by the fastest open-source model, DeepSeek.
For Usability, GPT-4o and GPT-4o-Mini lead in all metrics with similar performance, while DeepSeek closely follows and performs similarly to GPT-3.5. Additionally, as shown in Figure 2, GPT-4o-Mini and DeepSeek exhibit the most diverse AST node types, indicating their mutations are more expressive.
For Behavior, GPT-4o and GPT-4o-Mini excel in all metrics, and DeepSeek has similar performance to GPT-3.5, making it the best open-source LLM. StarChat and CodeLlama underperform across most metrics, particularly in compilation rate and bug detection. Based on the model information in Table 2, we find that newer models tend to generate higher-quality mutations, which aligns with the trends observed in other code-related tasks (Yang et al., 2024).
Finding 7 (RQ3): Two GPT-4 models perform best, while DeepSeek has a similar performance to GPT-3.5, standing out among open-source models. The newer models exhibit better performance in mutation generation.
4.4. RQ4: Impacts of Different Prompts
Table 7. Comparison Results of Different Prompts
Metric Category Metric GPT-3.5 DC-236b
P1 P2 P3 P4 P1 P2 P3 P4
—— Mut. Score 0.708 0.706 0.736 0.782 0.712 0.723 0.719 0.737
Mut. Count 48670 49126 42775 40844 53000 53177 52687 53130
Cost Avg. Gen. Time 1.79 1.78 1.80 1.87 4.25 4.25 4.20 4.40
Usability Compilability Rate 65.7% 63.8% 61.5% 60.6% 74.2% 75.1% 70.2% 72.6%
Useless Mut. Ratio 13.6% 11.3% 21.8% 14.8% 8.9% 3.8% 11.7% 9.8%
Eq. Mut. Rate 2.2% 1.9% 2.4% 2.5% 1.2% 1.3% 1.0% 1.5%
Behavior Real Bug Detectability 81.4% 80.7% 75.5% 68.0% 85.0% 83.7% 82.0% 82.7%
Coupling Rate 15.4% 15.1% 11.6% 11.7% 14.8% 14.6% 13.9% 14.4%
Ochiai Coefficient 66.3% 65.4% 63.1% 64.7% 66.7% 64.1% 64.4% 65.0%
Table 7 presents the comparative results of GPT-3.5 and DeepSeek using different prompts listed in Section 3.5.3. Prompts P1 to P3 simplify progressively, each containing less information than its predecessor, while P4 is the most complex, enhancing P1 with test suite code.
Overall, P1, the default prompt, excels in all Behavior Metrics on both models, indicating it can generate mutations more likely to resemble real bugs and effectively detect them. In addition, for all Usability Metrics, P1 achieves the second-best performance. P2, created by removing few-shot examples from P1, performs better in usability. P3, provided only with the code element to be mutated, achieves the lowest equivalent mutation rate. Conversely, P4, which extends P1 with additional test suite code, performs the worst in most cases.
Finding 8 (RQ4): The prompt with the whole method and few-shot examples as context (i.e., P1) achieves the best performance across all Behavior Metrics, whereas adding the code of test suites (i.e., P4) decreases performance.
4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 4.1, LLMs generate a significant number of non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT-3.5, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified 9 distinct error types, as shown in Table 8.
@@ 4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Shown as Table 8, the most common error type, Usage of Unknown Methods, accounts for 27.3% of the total errors, revealing the hallucination issue of generative models (Huang et al., 2023). Code Structural Destruction is the second most common error, accounting for 22.9%, indicating that ensuring generated code is syntactically correct remains a challenge for current LLMs. This result suggests that there is still significant room for improvement in current LLMs.
Finding 9 (RQ5): The mutations generated by GPT models have nine main types of compilation errors, with Usage of Unknown Methods and Code Structural Destruction being the two most prevalent types.
(a) GPT-3.5
(b) GPT-4o
(c) GPT-4o-M
(d) DS-236b
(e) SC-16b
(f) CL-13b
(g) LEAM
(h) $\mu$Bert
(i) Major
@@
Finding 10 (RQ5): Member references and method invocations within the code context are the most likely triggers for LLMs to generate non-compilable mutations.
@@ 5. Discussion
5.1. Sensitivity to Chosen Experiment Settings
@@ 5.1.1. Experiment Setup of the Context Length
Metric Category Metric 1-Line Context 2-Line Context 3-Line Context
Usability Compilability Rate 0.610 0.617 0.657
Useless Mutation Ratio 0.119 0.128 0.137
Equivalent Mutation Rate 1.6% 1.1% 1.3%
Behavior Real Bug Detectability 0.933 0.917 0.917
Coupling Rate 0.141 0.121 0.136
Ochiai Coefficient 0.517 0.417 0.467
Table 10. Performance Similarity of Different Context Lengths
@@
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 9. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the upper part of Table 10. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.

### v3

@@ 3.4.1. Categorizations of compilation error types
    Usage of Unknown Methods: this type of error is due to invoking unknown methods under the code context.
@@
    Code Structural Destruction: this type of error is due to illegal code structures, such as unmatched parentheses or braces.
@@
    Incorrect Method Parameters: this type of error is due to invoking a method with illegal parameters, including invoking by wrong numbers of parameters or unmatched type.
@@
    Usage of Unknown Variables: this type of error is due to using unknown variables under the code context.
  5. (5)
    Usage of Unknown Types: this type of error is due to using unknown type names under the code context.
  6. (6)
    Type Mismatch: this type of error is due to using illegal types when applying a certain operation, such as using objects to perform arithmetic operations.
  7. (7)
    Incorrect Initialization: this type of error is due to using illegal forms of constrictors or deleting necessary initialization statements.
  8. (8)
    Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
  9. (9)
    Incorrect Exceptions: this type of error is due to throwing illegal exceptions.
@@
Table 11 presents the results. From the table, we observe that the most frequent error type, Usage of Unknown Methods, accounts for 27.3% of total errors, highlighting the challenge of hallucination in LLM-generated code (Huang et al., 2023). The second most common error, Code Structural Destruction, makes up 22.9%, indicating that maintaining syntactic correctness remains a persistent difficulty for LLMs. These findings underscore the need for further improvements in LLMs to enhance their reliability and accuracy in code generation.
(a) Major
(b) LEAM
(c) $\mu$Bert
(d) GPT-3.5
(e) GPT-4o
(f) GPT-4o-M
(g) DS-236b
(h) SC-16b
(i) CL-13b
@@ 3.4.2. Categorizations of code context of non-compilable mutations
Answer to RQ4: The mutations generated by GPTs have nine main types of compilation errors, with Usage of Unknown Methods and Code Structural Destruction being the most prevalent. Deletion mutations are not the primary cause of non-compilable code. An analysis of the surrounding code context where GPTs produce non-compilable mutations reveals that member references and method invocations are the most frequently occurring code features associated with these errors.
@@ 4. Discussion
In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
4.1. Sensitivity to Experimental Settings
We discuss the experiment setups that may affect the evaluation results, particularly the setups related to LLMs.
4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutation generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get four types of prompts, shown as follows:
  1. (1)
  2. (2)
    Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
  3. (3)
  4. (4)
Table 12. Comparative Evaluation of Different Prompts for LLMs
Metric GPT-3.5 DS-236b
P1 P2 P3 P4 P1 P2 P3 P4
Mut. # 3681 3688 3350 3210 3803 3845 3802 3804
Mut. Score 0.773 0.785 0.842 0.801 0.784 0.778 0.793 0.814
Real Bug Detec. 79.4% 78.4% 72.2% 72.5% 84.6% 84.0% 81.4% 82.6%
Avg. Ochiai 63.1% 61.8% 60.8% 62.1% 67.3% 64.1% 65.0% 63.7%
Coupling Rate 52.1% 52.2% 47.4% 49.6% 53.1% 51.9% 52.9% 50.7%
Compilability Rate 66.9% 66.8% 60.0% 61.2% 79.1% 78.9% 74.9% 78.2%
Dup. Mut. Rate 11.4% 13.3% 14.9% 12.7% 10.9% 9.4% 9.8% 12.7%
Avg. Gen. Time 1.789 1.782 1.809 1.884 4.252 4.254 4.199 4.397
Table 12 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek across multiple metrics. Among the four prompts, P1 consistently demonstrates superior performance across both models, achieving the highest Compilability Rates (66.9% for GPT-3.5 and 79.1% for DeepSeek) and the best Real Bug Detectability Rates (79.4% and 84.6%, respectively). P1 also yields the highest Avgerage Ochiai Coefficient across both models. These results overall indicate that providing LLMs with well-structured and appropriate information significantly enhances their performance (e.g., P1), whereas excessive information can lead to a decline in effectiveness (e.g., P4).
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 14. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05.
@@ 4.1.2. Experiment Setup of the Context Length
Metric 1-Line Context 2-Line Context 3-Line Context
Real Bug Dectc. 85.0% 83.3% 83.3%
Avg. Ochiai 33.3% 35.0% 30.0%
Coupling Rate 36.6% 35.2% 37.1%
Compilability Rate 62.9% 63.2% 67.2%
Dup. Mut. Rate 10.8% 12.5% 11.9%
Table 14. Performance Similarity of Different LLM Configurations
@@
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 13. Similarly, the middle part of Table 14 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95 and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.